The photochemical formation of molecule Z uses a radiation of 300 nm. The sample is  irradiated with 100W of this radiation. The quantum yield is 0.2. If the sample absorbs only  50% of the radiation falling on it, then find the amount of Z formed in one second


1
Expert's answer
2021-01-06T08:31:44-0500

Photon energy is E = hc/l = 6.626*10-34*3*108/3*10-7 = 6.626*10-19 J

Power is energy per second, hence, number of photons per second is:

N(photon) = 100/6.626*10-19 = 1.509*1020

The amount of Z formed per second is:

N(Z) = N(photon)*0.5(absorption)*0.2(quantum yield) = 1.509*1019 molecules per second, or,

1.509*1019/6.02*1023 = 2.507*10-5 mole per second.